OpenStudy (anonymous):

Write the equation for the circle that satisfies the following conditions: Center (-11,-10) and radius 8 units. a. (x+11)^2 + (y+10)^2 = 64 b. (x-11)^2 + (y+10)^2 = 64 c. (x+11)^2 + (y-10)^2 = 64 d. (x+11)^2 + (y+10)^2 = 8 How do I find which answer is right?

Directrix (directrix):

You need to know the standard form for the equation of a circle so that you can identify the center and radius. (x -( h)) ^2 + (y -( k))^2 = r^2 h = -11, k = -10, and r = 8 ( x - (-11) ) ^2 + ( y - ( -10 ) ) ^2 = 8^2 (x + 11) ^2 + ( y + 10 ) ^2 = 64 a. (x+11)^2 + (y+10)^2 = 64 ---- Option A